Joe is 4 times as old as Bob. Bob is 5 years older than Phil. All their ages sum to 73. What is the Difference between Joe and Phil's ages?
Answer by TimothyLamb(4379)

---
x = joe
y = bob
z = phil
---
x = 4y
y = z + 5
x + y + z = 73
---
put the system of linear equations into standard form
---
x - 4y = 0
y - z = 5
x + y + z = 73
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x= joe = 52
y= bob = 13
z= phil = 8
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php